Ship reports banner with links Home IODP-USIO daily/weekly reports

< January

IODP Expedition 345

Hess Deep Plutonic Crust

Daily Science Report for 11 February 2013

LOCATION: In transit to Balboa, Panama

SCIENCE UPDATE: We held our final meeting to discuss postcruise research and publications. All shipboard science reports and activities are complete. As of 2400 h on 11 February, we have transited a total of 1397 nmi since leaving Hess Deep for Balboa, Panama. We have 7 nmi remaining and expect to be dockside in Balboa at ~1200 h on 12 February 2013.


Daily Science Report for 10 February 2013

LOCATION: In transit to Balboa, Panama

SCIENCE UPDATE: As of 0000 h on 11 February, we have transited a total of 1223 nmi since leaving Hess Deep for Balboa, Panama. We have 244 nmi remaining and expect to be dockside in Balboa at ~1200 h on 12 February 2013. We had our second meeting to discuss initial Hole U1415P results and scientists have completed all site reports.


Daily Science Report for 9 February 2013

LOCATION: In transit to Balboa, Panama

SCIENCE UPDATE: Our transit from Hess Deep to Balboa, Panama continues. All samples from Expedition 345 cores have been taken. A few scientists finalized selection of ODP Leg 147 cores for postcruise research. A meeting was held to discuss initial Hole U1415P results and scientists spent the rest of the day working on their final site reports. We are scheduled to arrive in Balboa on 12 February.


Daily Science Report for 8 February 2013

LOCATION: In transit to Balboa, Panama

SCIENCE UPDATE: Today, we continued our transit from Hess Deep to Balboa, Panama. Scientists completed selection of samples for postcruise research and working on their final site reports. We are scheduled to arrive in Balboa on 12 February.


Daily Science Report for 7 February 2013

LOCATION: In transit to Balboa, Panama

SCIENCE UPDATE: Our transit from Hess Deep to Balboa, Panama continued. The first samples for postcruise research were cut and adjustments were made to samples in areas of high demand. Scientists continued working on their final reports. We are scheduled to arrive in Balboa on 12 February.


Daily Science Report for 6 February 2013

LOCATION: In transit to Balboa, Panama

SCIENCE UPDATE: Today was our first full day of transit from Hess Deep to Balboa, Panama. Scientists completed the final laboratory analyses, prepared their research and core sampling plans, and worked on their final reports. We are scheduled to arrive in Balboa on 12 February.


Daily Science Report for 5 February 2013

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: We retrieved the logging tools back on the rig floor at 0030 h. Since our logging tools were not able to pass below 24 mbsf, we tried lowering the end of the pipe past this depth. However, the pipe could not pass below this depth even with 10,000 lb weight and while circulating. So, we abandoned our attempts to log. We pulled the pipe out of Hole U1415P and conducted a short camera and 3.5 kHz subbottom survey across Hole U1415P. After the survey was completed at 0845 h, we started recovering the camera and drill sting. The end of the pipe was back on the rig floor at 1745 h. Since our final run of the ICP for geochemical analyses was still running, we delayed our departure for Panama until 2045 h.


Daily Science Report for 4 February 2013

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: We spent the first half of the day attempting to stabilize problematic hole conditions between 79 mbsf and the bottom of the hole at 107.9 mbsf. Given the short time remaining in the expedition and our interest in trying to acquire downhole log data, we decided to cease attempts to clean out the bottom of the hole. To prepare for logging, we raised the bit up to 12 mbsf into the bottom of the 10.75 inch casing and lowered it back down the hole. Several problematic intervals had to be drilled out to advance down the hole. We did not attempt to get past 90 mbsf. We pulled the bit out of the hole, offset the ship, released the bit on the seafloor, reentered the hole, and placed the end of the pipe at 10 mbsf. At 1825 h, we assembled the first logging string and lowered it to the end of the pipe, but it would not pass 24 mbsf (11.5 m below the casing). At the end of the day, we were retrieving the logging tools so that we could try lowering the pipe below the obstruction.


Daily Science Report for 3 February 2013

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: Today we cut Cores U1415P-20R, ‑22R, and ‑23R from 82.3 to 107.9 mbsf. These recovered 7.29 m of medium to coarse-grained gabbroic rocks (28%). We encountered problematic hole conditions below ~79 mbsf and spent nearly half the day conducting hole cleaning operations between cores.


Daily Science Report for 2 February 2013

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: After installing a new RCB bit, we reentered Hole U1415P just before 0600 h. We recovered the camera system and then spent from 0815 h to 1815 h washing and reaming back to the bottom of the hole at 82.3 mbsf. The last ~12 m took quite a while to penetrate. We recovered the core barrel from this washing and reaming at 1920 h (Core U1415P-19G; 0.60 m of gabbroic rocks) and resumed RCB coring.


Daily Science Report for 1 February 2013

LOCATION: Site U1415 (HD-01B), Hess Deep

SCIENCE UPDATE: Today, we RCB cored continuously from 73.0 to 82.3 mbsf. Cores U1415P-17R and ‑18R obtained 1.30 m of medium-coarse grained gabbroic rocks from this 9.3 m interval (14% recovery). After having some challenges drilling back to the bottom of the hole after each of the last two cores, we decided to retrieve the drill string to install a new RCB bit. Our plan, depending on hole conditions, is to core and then log until we depart for Panama at 1000 h on 6 February.




About the IODP JRSO | Expeditions | Data & core samples | Tools & laboratories | Publications | Education |
Newsroom | Meetings & port calls | Employment | Contact us | Search | Site map | People | Intranet | Home

Accesibility Statement | Plug-in download | Privacy statement | ODP

For comments or questions: webmaster@iodp.tamu.edu

Copyright IODP JRSO

Modified on Friday, 01-Mar-2013 10:47:28 CST.